OpenStudy (anonymous):

Solve for the variable. y - 10 = 14

OpenStudy (carson889):

Add 10 to both sides. y - 10 + 10 = 14 + 10 y = 24

hartnn (hartnn):

add 10 on both sides, what u get ?

OpenStudy (anonymous):

y-10=14, y=14+10, y=24. Answer: 24.